# Neural Dueling Bandits

### v1

@@ 3.2 Regret analysis
As we have discussed above, compared with the regret upper bounds of NeuralUCB [7] and NeuralTS [8], the effective dimension $\widetilde{d}$ in 2 and 3 are expected to be larger than the effective dimension $\widetilde{d}^{\prime}$ from [7, 8] because our $\widetilde{d}$ results from the summation of a significantly larger number of contexts. Therefore, our regret upper bounds (2 and 3) are expected to be worse than that of [7, 8]: $\widetilde{O}(\widetilde{d}^{\prime}\sqrt{T})$. This downside can be attributed to the difficulty of our neural dueling bandits setting, in which we can only access preference feedback.
4 Theoretical Insights for Reinforcement Learning with Human Feedback
Our algorithms and theoretical results can also provide insights on the celebrated reinforcement learning with human feedback (RLHF) algorithm [39], which has been the most widely used method for the alignment of large language models (LLMs). In RLHF, we are given a dataset of user preferences, in which every data point consists of a prompt and a pair of responses generated by the LLM, as well as a binary observation indicating which response is preferred by the user. Following our notations in Section 2, the action $x_{t,1}$ (resp. $x_{t,2}$) corresponds to the concatenation of the prompt and the first response (resp. second response). Of note, RLHF is also based on the assumption that the user preference over a pair of responses is governed by the BTL model (Section 2). That is, the binary observation $y_{t}$ is sampled from a Bernoulli distribution, in which the probability that the first response is preferred over the second response is given by $\mathbb{P}\left\{x_{t,1}\succ x_{t,2}\right\}=\mu\left(f(x_{t,1})-f(x_{t,2})\right)$. Here $f$ is often referred to as the reward function, which is equivalent to the latent utility function $f$ in our setting (Section 2).
Typically, RLHF consists of two steps: (a) learning a reward model using the dataset of user preferences and (b) fine-tuning the LLM to maximize the learned reward model using reinforcement learning. In step (a), same as our algorithms, RLHF also uses an NN $h$ (which takes as input the embedding from a pre-trained LLM) to learn the reward model by minimizing the loss function (1). The accuracy of the learned reward model is crucial for the success of RLHF [39]. Importantly, our 1 provides a theoretical guarantee on the quality of the learned reward model $h$, i.e., an upper bound on the estimation error of the estimated reward differences between any pair of responses. Therefore, our 1 provides a theoretically principled measure of the accuracy of the learned reward model in RLHF, which can potentially be used to evaluate the quality of the learned reward model.
In addition, some recent works have proposed the paradigm of online/iterative RLHF [40, 41, 42, 43, 44] to further improve the alignment of LLMs. In online RLHF, the RLHF procedure is repeated multiple times. Specifically, after an LLM is fine-tuned to maximize the learned reward model, it is then used to generate pairs of responses to be used to query the user for preference feedback; then, the newly collected preference data is added to the preference dataset to be used to train a new reward model, which is again used to fine-tune the LLM. In this case, as the alignment of the LLM is improved after every round, the newly generated responses by the improved LLM are expected to achieve progressively higher reward values, which have been shown to lead to better alignment of LLMs [40, 41, 42, 43, 44]. In every round, we can let the LLM generate a large number of responses (i.e., the actions in our setting, see Section 2), from which we can use our algorithms to select two responses $x_{t,1}$ and $x_{t,2}$ to be shown to the user for preference feedback. In addition, our algorithm can also potentially be used to select the prompts shown to the user, which correspond to the contexts in our problem setting (Section 2). Our theoretical results guarantee that our algorithms can help select responses with high reward values (2 and 3). Therefore, our algorithms can be used to improve the efficiency of online RLHF.
5 Neural Contextual Bandits with Binary Feedback
We now extend our results to the neural contextual bandit problem in which a learner only observes binary feedback for the selected arms (note that the learner only selects one arm in every iteration). Observing binary feedback is very common in many real-life applications, e.g., click or not in online recommendation and treatment working or not in clinical trials [19, 45, 46].
Contextual bandits with binary feedback.
Performance measure.
5.1 Reward function estimation using neural network and our algorithms
UCB-based algorithm.
1:  Tuning parameters: $\delta\in(0,1)$ and $\lambda>0$
2:  for $t=1,\ldots,T$ do
6:     Observe preference feedback binary $y_{t}$
7:  end for
TS-based algorithm.
Regret analysis.
We use the above fact to prove the following confidence ellipsoid result as done in linear reward function [19, 45, 46].
Theorem 4.
for all $x\in\mathcal{X}_{t},t\in[T]$.
Theorem 5 (NCBF-UCB).
@@ 5.1 Reward function estimation using neural network and our algorithms
$\mathfrak{R}_{T}\leq 2\sqrt{2}\left(\beta_{T}+B\sqrt{\frac{\lambda}{\kappa_{\mu}}}+1\right)\sqrt{c_{0}T\widetilde{d}_{b}}+1=\widetilde{O}\left(\left(\frac{\sqrt{\widetilde{d}_{b}}}{\kappa_{\mu}}+B\sqrt{\frac{\lambda}{\kappa_{\mu}}}\right)\sqrt{T\widetilde{d}_{b}}\right)$
Theorem 6 (NCBF-TS).
Note that in terms of asymptotic dependencies (ignoring the log factors), our UCB- and TS- algorithms have similar growth rates. All missing proofs and additional details are in Appendix C.
Comparison with Neural Bandits. The regret upper bounds of our NCBF-UCB and NCBF-TS algorithms are worse than the regret of NeuralUCB [7] and NeuralTS [8]: $\widetilde{O}(\widetilde{d}_{b}\sqrt{T})$ (with $\kappa_{\mu}=1$) because of our extra dependency on $\kappa_{\mu}$ and $L_{\mu}$. [^4] Note that our effective dimension $\widetilde{d}_{b}$ is defined using $\mathbf{H}_{b}$ Eq. 8, while the effective dimension $\widetilde{d}^{\prime}$ in [7] and [8] are defined using $\mathbf{H}$. However, as we have discussed in Footnote 3, $\widetilde{d}^{\prime}$ has the same order of growth as $\log\det\left(\mathbf{H}_{b}/\lambda+\mathbf{I}\right)$. So, our regret upper bounds are comparable with those from [7] and [8]. Specifically, note that $\kappa_{\mu}<1$, therefore, the regret bounds in 5 and 6 are increased as a result of the dependency on $\kappa_{\mu}$. In addition, the dependency on $L_{\mu}$ also places an extra requirement on the width $m$ of the NN. Therefore, our regret bounds are worse than that of standard neural bandit algorithms that do not depend on $\kappa_{\mu}$ and $L_{\mu}$. This can be attributed to the additional difficulty of our problem setting, i.e., we only have access to binary feedback, whereas standard neural bandits [7, 8] can use continuous observations.
@@ 6 Experiments
To corroborate our theoretical results, we empirically demonstrate the performance of our algorithms on different synthetic reward functions. We adopt the following two synthetic functions from earlier work on neural bandits [7, 8, 15]: $f(x)=10(x^{\top}\theta)^{2}$ (Square) and $f(x)=\cos(3x^{\top}\theta)$ (Cosine). We repeat all our experiments 20 times and show the average and weak cumulative regret with a 95% confidence interval (represented by the vertical line on each curve) Due to space constraints, more details of the following experiments and additional results are given in the Appendix.
Regret comparison with baselines.  We compare regret (average/weak of our proposed algorithms with three baselines: LinDB-UCB (adapted from [1]), LinDB-TS, and CoLSTIM [2]. LinDB-UCB and LinDB-TS can be treated as variants of NDB-UCB and NDB-TS, respectively, in which a linear function approximates the reward function. As expected, NDB-UCB and NDB-TS outperform all linear baselines as these algorithms cannot estimate the non-linear reward function and hence incur linear regret. For a fair comparison, we used the best hyperparameters of LinDB-UCB and LinDB-TS (see Fig. 3) for NDB-UCB and NDB-TS. We observe the same trend for different non-linear reward functions (see Fig. 5 and Fig. 6) and also for NCBF-UCB and NCBF-TS (see Fig. 7).
@@
Varying dimension and arms vs. regret  Increasing the number of arms ($K$) and the dimension of the context-arm feature vectors $(d)$ makes the problem more difficult. To see how increasing $K$ and $d$ affects the regret of our proposed algorithms, we vary the $K=\{5,10,15,20,25\}$ and $d=\{5,10,15,20,25\}$ while keeping the other problem parameters constant. As expected, the regret of NDB-UCB increases with increase in $K$ and $d$ as shown in Fig. 2. We also observe the same behavior for NDB-TS as shown Fig. 4. All missing figures from this section are in the Appendix D.

### v2

@@ 3.3 Regret analysis
Compared with the regret upper bounds of NeuralUCB Zhou et al. (2020) and NeuralTS Zhang et al. (2021), the effective dimension $\widetilde{d}$ in 2 and 3 are expected to be larger than the effective dimension $\widetilde{d}^{\prime}$ from Zhou et al. (2020); Zhang et al. (2021) because our $\widetilde{d}$ results from the summation of a significantly larger number of contexts. Therefore, our regret upper bounds (2 and 3) are expected to be worse than that of Zhou et al. (2020); Zhang et al. (2021): $\widetilde{O}(\widetilde{d}^{\prime}\sqrt{T})$. This downside can be attributed to the difficulty of our neural dueling bandits setting, in which we can only access preference feedback.
3.4 Proof Sketch
In this section, we give a brief proof sketch of our regret analysis in Section 3.3.
3.4.1 Sketch of Proof for 1
We start by presenting an outline for the proof of our 1. To begin with, by applying the theory of the NTK (Jacot et al., 2018) to our NN for reward estimation (Section 3.1), we show that as long as the NN is wide enough, its output can be approximated by a linear function (see Section A.1 for details). More specifically, for a pair of arms $x$ and $x^{\prime}$, the difference between their predicted latent reward values $h(x;\theta_{t})-h(x^{\prime};\theta_{t})$ can be approximated by the difference between the linear approximations of the NN:
Lemma 1.
Let $\varepsilon^{\prime}_{m,t}\triangleq C_{2}m^{-1/6}\sqrt{\log m}L^{3}\left(\frac{t}{\lambda}\right)^{4/3}$ where $C_{2}>0$ is an absolute constant. Then
$|\langle\varphi(x)-\varphi(x^{\prime}),\theta_{t}-\theta_{0}\rangle-(h(x;\theta_{t})-h(x^{\prime};\theta_{t}))|\leq 2\varepsilon^{\prime}_{m,t},\,\,\,\forall t\in[T],x,x^{\prime}\in\mathcal{X}_{t}.$
Of note, the approximation error $\varepsilon^{\prime}_{m,t}$ becomes smaller as the width $m$ of the NN is increased.
Proof of the Confidence Ellipsoid.
Next, we derive the confidence ellipsoid for our algorithms, which is formalized by 6 (see its detailed proof in Section A.2). 6 allows us to derive the following inequality, which shows that for a pair of arms $x$ and $x^{\prime}$, the difference between their latent reward values $f(x)-f(x^{\prime})$ can be approximated by the difference between their predicted values by the linearized NN:
@@ 3.4.1 Sketch of Proof for 1
$|f(x)-f(x^{\prime})-\langle\varphi(x)-\varphi(x^{\prime}),\theta_{t}-\theta_{0}\rangle|\leq\left\|\frac{1}{\sqrt{m}}\left(\varphi(x)-\varphi(x^{\prime})\right)\right\|_{V_{t-1}^{-1}}\left(\beta_{T}+B\sqrt{\frac{\lambda}{\kappa_{\mu}}}+1\right),$ (5)
Importantly, Eq. 5 guarantees that our trained NN with parameters $\theta_{t}$ (after linearization) is an accurate approximation of the latent reward function $f$ in terms of pairwise differences. This crucial theoretical guarantee is achieved thanks to our carefully designed loss function Eq. 1. Specifically, a key step in the proof of our confidence ellipsoid (i.e., 6) is the following equality:
$\frac{1}{m}\sum_{s=1}^{t-1}\left(\mu\left(h(x_{s,1};\theta_{t})-h(x_{s,2};\theta_{t})\right)-y_{s}\right)(g(x_{s,1};\theta_{t})-g(x_{s,2};\theta_{t}))+\lambda(\theta_{t}-\theta_{0})=0,$ (6)
which is achieved by obtaining the NN parameters $\theta_{t}$ by minimizing our loss function Eq. 1. This is, in fact, analogous to using maximum likelihood to estimate the unknown vector $\theta$ in the linear reward function in linear dueling bandits (Bengs et al., 2022) and linear bandits with binary feedback (Li et al., 2017). In addition, another important step in the proof of 6 is the recognition that the observation noise $\varepsilon_{t}=y_{t}-\mu(f(x_{t,1})-f(x_{t,2}))$ is conditionally $1$-sub-Gaussian (more details can be found in the proof of 8 in Section A.2). After the proof of the confidence ellipsoid as described above, we then combine Eq. 5 and 1 to derive an upper bound on the difference between $f(x)-f(x^{\prime})$ and $h(x;\theta_{t})-h(x^{\prime};\theta_{t})$, which completes the proof of 1.
3.4.2 Sketch of Proof for 2 and 3
To obtain a regret upper bound for the NDB-UCB algorithm (i.e., to prove 2), we firstly use 1 to derive an upper bound on the average instantaneous regret $r_{t}^{a}=f(x_{t}^{\star})-\left({f(x_{t,1})+f(x_{t,2})}\right)/{2}$ (see the detailed proof in Section A.3):
$r_{t}^{a}\leq\frac{3}{2}\left(\beta_{T}+B\sqrt{\frac{\lambda}{\kappa_{\mu}}}+1\right)\sqrt{\frac{\kappa_{\mu}}{\lambda}}\sigma_{t-1}(x_{t,1},x_{t,2})+3\varepsilon^{\prime}_{m,t}.$ (7)
Our strategies to select the two arms $x_{t,1}$ and $x_{t,2}$ (i.e., Eq. 2 and Eq. 3) are essential for deriving the upper bound in Eq. 7. Subsequently, we follow similar steps to the analysis of the kernelized bandit algorithms (e.g., GP-UCB (Srinivas et al., 2010)) to obtain an upper bound on $\sum^{T}_{t=1}\sigma_{t-1}(x_{t,1},x_{t,2})$ in terms of $\widetilde{d}$. This then provides us with an upper bound on the average regret $\mathfrak{R}_{T}^{a}=\sum^{T}_{t=1}r^{a}_{t}$ (as well as the weak regret $\mathfrak{R}_{T}^{w}$ since $\mathfrak{R}_{T}^{w}\leq\mathfrak{R}_{T}^{a}$) and hence completes the proof of 2. The proof of 3 for the NDB-TS algorithm is also based on 1 and largely follows from the analysis of the GP-TS algorithm (Chowdhury & Gopalan, 2017). However, non-trivial modifications need to be made in order to incorporate our strategies to select the two arms $x_{t,1}$ and $x_{t,2}$. The detailed regret analysis for 2 and 3 can be found in Section A.3 and Appendix B, respectively.
4 Further Implications of Our Algorithms and Theory
Our results to learn the latent non-linear reward function using preference feedback is not only limited to dueling bandits, but it can be adapted to other settings as discussed in the following sections.
4.1 Neural Contextual Bandits with Binary Feedback
In many real-world applications, the learner can only observe binary feedback, e.g., whether a user clicks in an online recommendation system or whether a treatment is effective in clinical trials, and these problems are commonly modeled as contextual GLM bandits (Filippi et al., 2010; Li et al., 2017; Faury et al., 2020), which assume the probability of success (i.e., $1$) is proportional to exponential of its reward that is a linear function of action-context features. However, in many real-life applications, the reward can be a non-linear function of action-context features, which can modeled using a suitable neural network. We refer to this problem as neural contextual bandits with binary feedback. Therefore, our results can be extended to the neural contextual bandit problem, where the learner only receives binary feedback for the selected arms, which depends on the non-linear reward function. This adaption is not straightforward as the learner selects only one arm and observes binary feedback (success or failure) for the selected action. This difference leads to different arm selection strategies as only one needs to be selected in neural contextual bandits, and different loss functions must be optimized to train the neural network to estimate the unknown reward functions. The main challenge is to derive the confidence ellipsoid result, which is a crucial component for proving the regret bounds. While some ideas are common, the derivations differ due to the underlying differences in the problem settings. We have given more details of the neural contextual bandit with binary feedback problem and our regret bounds of algorithms proposed for this setting in Appendix D.
4.2 Reinforcement Learning from Human Feedback
Our algorithms and theoretical results can also provide insights on the celebrated reinforcement learning with human feedback (RLHF) algorithm (Chaudhari et al., 2024), which has been the most widely used method for aligning large language models (LLMs). Firstly, our Theorem 1 provides a theoretical guarantee on the quality of the learned reward model during RLHF. This is because Theorem 1 serves as an upper bound on the estimation error of the estimated reward differences between any pair of prompts. Secondly, our proposed algorithms can be used to improve the quality of the human preference dataset for RLHF. That is when collecting the dataset for RLHF, we can let the LLM generate a large number of responses, from which we can use our algorithms for neural dueling bandits to select two responses (corresponding to two arms) to be shown to the user for preference feedback. More details can be found in Appendix E.
@@ 5 Experiments
To corroborate our theoretical results, we empirically demonstrate the performance of our proposed algorithms on different synthetic reward functions. We adopt the following two synthetic functions from earlier work on neural bandits (Zhou et al., 2020; Zhang et al., 2021; Dai et al., 2023): $f(x)=10(x^{\top}\theta)^{2}$ (Square) and $f(x)=\cos(3x^{\top}\theta)$ (Cosine). We repeat all our experiments 20 times and show the average and weak cumulative regret with a 95% confidence interval (represented by the vertical line on each curve) Due to space constraints, additional results are given in the Appendix C.
The details of our experiments are as follows: We use a $d$-dimensional space to generate the sample features of each context-arm pair. We denote the context-arm feature vector for context $c_{t}$ and arm $a$ by $x_{t}^{a}$, where $x_{t}^{a}=\left(x_{t,a}^{1},\ldots,x_{t,a}^{d}\right),~{}\forall~{}t\geq 1$. The value of $i$-the element of $x_{t}^{a}$ vector is sampled uniformly at random from $\left(-1,1\right)$. Note that the number of arms remains the same across the rounds, i.e., $K$. We then select a $d$-dimensional vector $\theta$ by sampling uniformly at random from $(-1,1)^{d}$. In all our experiments, the binary preference feedback about $x_{1}$ being preferred over $x_{2}$ (or binary feedback in Appendix D) is sampled from a Bernoulli distribution with parameter $p=\mu\left(f(x_{1})-f(x_{2})\right)$). In all our experiments, we use a NN with $2$ hidden layers with width $50$, $\lambda=1.0$, $\delta=0.05$, $d=5$, $K=5$, and fixed value of $\nu_{T}=\nu=1.0$. For having a fair comparison, we choose the value of $\nu$ after doing a hyperparameter search over set $\{10.0,5.0,1.0,0.1,0.01,0.001,0.0\}$ for linear baselines, i.e., LinDB-UCB and LinDB-TS. As shown in Fig. 1, the average cumulative regret is minimum for $\nu=1.0$. Note that we did not perform any hyperparameter search for NDB-UCB and NDB-TS, whose performance can be further improved by doing the hyperparameter search.
(a) Square (UCB)
(b) Square (TS)
(c) Cosine (UCB)
(d) Cosine (TS)
As supported by the neural tangent kernel (NTK) theory, we can substitute the initial gradient $g(x;\theta_{0})$ for the original feature vector $x$ as $g(x;\theta_{0})$ represents the random Fourier features for the NTK Jacot et al. (2018). In this paper, we use the feature vectors $g(x;\theta_{t})$ instead of $g(x;\theta_{0})$ and recompute all $g(x;\theta_{t})$ in each round for all past context-arm pairs. Additionally, compared to NTK theory, we have designed our algorithm to be more practical by adhering to the common practices in neural bandits Zhou et al. (2020); Zhang et al. (2021). Specifically, in the loss function for training our NN (as defined in Eq. 1), we replaced the theoretical regularization parameter $\frac{1}{2}m\lambda\left\|\theta-\theta_{0}\right\|^{2}_{2}$ (where $m$ is the width of the NN) with the simpler $\lambda\left\|\theta\right\|^{2}_{2}$. Similarly, for the random features of the NTK, we replaced the theoretical $\frac{1}{\sqrt{m}}g(x;\theta_{t})$ with $g(x;\theta_{t})$. We retrain the NN after every $20$ rounds and set the number of gradient steps to $50$ in all our experiments.
Regret comparison with baselines.  We compare regret (average/weak of our proposed algorithms with three baselines: LinDB-UCB (adapted from (Saha, 2021)), LinDB-TS, and CoLSTIM (Bengs et al., 2022). LinDB-UCB and LinDB-TS can be treated as variants of NDB-UCB and NDB-TS, respectively, in which a linear function approximates the reward function. As expected, NDB-UCB and NDB-TS outperform all linear baselines as these algorithms cannot estimate the non-linear reward function and hence incur linear regret. For a fair comparison, we used the best hyperparameters of LinDB-UCB and LinDB-TS (see Fig. 1) for NDB-UCB and NDB-TS. We observe the same trend for different non-linear reward functions (see Fig. 5 and Fig. 6 in Appendix C) and also for neural contextual bandit with binary feedback (see Fig. 7 in Appendix D).
@@
Varying dimension and arms vs. regret  Increasing the number of arms ($K$) and the dimension of the context-arm feature vectors $(d)$ makes the problem more difficult. To see how increasing $K$ and $d$ affects the regret of our proposed algorithms, we vary the $K=\{5,10,15,20,25\}$ and $d=\{5,10,15,20,25\}$ while keeping the other problem parameters constant. As expected, the regret of NDB-UCB increases with increase in $K$ and $d$ as shown in Fig. 3. We also observe the same behavior for NDB-TS as shown Fig. 4. All missing figures from this section are in the Appendix C.
@@
(a) Varying $d$ (Average)
(b) Varying $d$ (Weak)
(c) Varying $K$ (Average)
(d) Varying $K$ (Weak)
Practical challenges.  Applying algorithms in scenarios requiring fast online interactions can be challenging for two reasons: updating the neural network (NN) and the arm-selection strategy. To address the first challenge, we can use a batched version of our algorithm, where the NN is updated only after a fixed interval. To address the second challenge, we can use an $\varepsilon_{t}$-greedy method for selecting arms when computing optimistic values using UCB or TS is not computationally feasible.
Computational resources used for all experiments.  All the experiments are run on a server with AMD EPYC 7543 32-Core Processor, 256GB RAM, and 8 GeForce RTX 3080.
6 Related Work
We now review the relevant work, especially in dueling bandits and contextual bandits, to our problem.
Finite-Armed Dueling Bandits.  Learning from pairwise or $K$-wise comparisons has been thoroughly explored in the literature. In the context of finite-armed dueling bandits, the learner only observes a pairwise preference between two selected arms, and the goal is to find the best arm (Yue & Joachims, 2009; 2011; Yue et al., 2012). In dueling bandit literature, different criteria (e.g., Borda winner, Condorcet winner, Copeland winner, or von Neumann winner) are used to find the best arm while focusing on minimizing the regret only using preference feedback (Zoghi et al., 2014b; Ailon et al., 2014; Zoghi et al., 2014a; Komiyama et al., 2015; Gajane et al., 2015; Saha & Gopalan, 2018; 2019a; 2019b; Zhu et al., 2023). We refer the readers to (Bengs et al., 2021) for a detailed survey on various works covering different settings of dueling bandits.
Contextual Bandits.  Many real-life applications in online recommendation, advertising, web search, and e-commerce can be modeled as contextual bandits (Slivkins, 2019; Lattimore & Szepesvári, 2020). In the contextual bandit setting, a learner receives a context (information before selecting an arm), selects an action for that context, and then receives a reward for the selected action. To deal with the large (or infinite) number of context-action pairs, the mean reward of each action is assumed to be parameterized by an unknown function of action-context features, e.g., linear (Li et al., 2010; Chu et al., 2011; Abbasi-Yadkori et al., 2011; Agrawal & Goyal, 2013b), GLM (Filippi et al., 2010; Li et al., 2017; Jun et al., 2017), and non-linear (Valko et al., 2013; Chowdhury & Gopalan, 2017; Zhou et al., 2020; Zhang et al., 2021). We adopt the simplest neural contextual bandit algorithms for dealing with non-linear rewards, i.e., NeuralUCB (Zhou et al., 2020) and NeuralTS (Zhang et al., 2021) to our setting, as also done in earlier works (Dai et al., 2023; Lin et al., 2023). Since NeuralUCB or NeuralTS primarily influences the arm selection strategy, we can incorporate any variants of these algorithms by making appropriate modifications to 2 and 1. These adoptions can be challenging because of our setting solely relies on pairwise comparisons.
Contextual Dueling Bandits.  The closest work to ours is contextual dueling bandits (Saha, 2021; Saha & Krishnamurthy, 2022; Bengs et al., 2022; Di et al., 2023; Li et al., 2024). Specifically, Saha (2021); Saha & Krishnamurthy (2022) proposed algorithms for contextual linear dueling bandits with pairwise and subset-wise preference feedback and established lower bounds for contextual preference bandits using a logistic link function. Whereas Bengs et al. (2022) generalized the setting to the contextual linear stochastic transitivity model, Di et al. (2023) considered the variance-aware algorithm, and Li et al. (2024) proposed an algorithm based on Feel-Good Thompson Sampling (Zhang, 2022). However, there are two key differences: non-linear reward function and arm-selection strategy. Existing work only considers the linear reward function, which may not be practical in many real-life applications. Our work fills this gap in the literature and generalizes the existing setting by considering the non-linear reward function in contextual dueling bandits. Furthermore, the existing works use different ways to select the pair of arms, leading to different arm-selection strategies than ours. Due to the differences in arm-selection strategy and non-linear reward function (which is estimated using an NN), our regret analysis is completely different than existing works.